# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### 12 Mar 2024 v1

@@ Abstract
With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that evaluates both spatial and textual relevance, have found many real-life applications. Existing geo-textual indexes for TkQs use traditional retrieval models like BM25 to compute text relevance and usually exploit a simple linear function to compute spatial relevance, but its effectiveness is limited. To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues. To the best of our knowledge, there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
@@
To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST). LIST is featured with two novel components. Firstly, we propose a lightweight and effective relevance model that is capable of learning both textual and spatial relevance. Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. Two key challenges in building an effective and efficient index are the absence of high-quality labels and unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
I INTRODUCTION
@@
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps [1]; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to retrieve geo-textual objects with texts relevant to the keywords ‘delicious Italian restaurants’ and located near the query’s location, and the result is a list of $k$ geo-textual objects ranked by a relevance model considering both spatial and textual relevance.
@@
Spatial keyword queries have applications in many scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions. To meet diverse user needs, various spatial keyword queries have been introduced [2, 3, 4]. Among them, the Top-k KNN Spatial Keyword Query (TkQ[2] retrieves top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. According to the experimental evaluation [5], TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query [4], which retrieves the geo-textual objects containing all the query keywords and rank them according to their spatial relevance (or proximity) to the query location. Most of the existing studies on spatial keyword queries focus on the improvement in the efficiency of handling spatial keyword queries. As such, many indexes [6] and corresponding query processing algorithms have been developed.
@@
Motivations. The existing indexes [2, 7, 8, 9] designed for the TkQ query typically compute textual relevance using traditional retrieval models such as BM25 [10] and TF-IDF [11], which rely on exact word matching to evaluate textual relevance. These traditional models often suffer from the word mismatch issue, i.e., synonyms or paraphrases that consist of completely different tokens may convey the same meanings, which limits their effectiveness (as to be shown in Section III-B). For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models, although they are relevant to the query.
As for spatial relevance, existing indexes [2, 7, 8, 9] for the TkQ query typically compute spatial relevance by a linear function of the distance between the query location and the location of a geo-textual object. The implicit assumption behind this practice is that the spatial preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we find it does not hold on our real-life datasets. From our dataset, we find that the spatial preference for geo-textual objects is not a linear function of distance, but exhibits a significant (non-linear) decrease with the increase of distance (detailed in Section III-B). This motivates us to design a more effective spatial relevance module. A straightforward approach would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (shown in Section V-E).
@@
To improve effectiveness, several deep learning based relevance models [5, 12, 13] propose to to compute textual relevance based on word embeddings. For example, DrW [5] employs embeddings generated from the Bert model, identifying top-k relevant terms from geo-textual object description for each query keyword based on the embeddings, and uses an attention mechanism to aggregate scores of each query keyword. Although these approaches could achieve better effectiveness, they have very high query latency for computing the relevance score since they rely on embeddings and neural networks (detailed in Section III-B). To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models.
Existing geo-textual indexes are developed based on traditional retrieval models such as TF-IDF. For instance, IR-Tree [2] augments each R-Tree [14] node with a pseudo-inverted file including the words of all objects inside that node. These indexes cannot be used to handle embedding based spatial keyword queries, which represent the semantic meaning of query keywords with a high-dimensional embedding. Although Approximate Nearest Neighbor Search (ANNS) indexes [15, 16, 17] are designed for embedding retrieval, none of these ANNS indexes take into account the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section V-B). It is an open problem to incorporate spatial factors into ANNS indexes to support embedding based spatial keyword queries.
Objectives. To address the aforementioned issues, this work aims to (1) design an embedding based lightweight yet effective spatio-textual relevance model as the ranking function, and (2) design a novel index to speed up high-dimensional embedding based spatial keyword queries.
To achieve the two objectives, we propose a novel solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries. We denote this retriever by LIST, which is featured with two new components. First, LIST develops a lightweight and effective relevance model as the ranking function, which comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module.
@@
Second, LIST introduces a novel machine learning based ANNS index. For both embedding based spatial keyword queries and geo-textual objects, our proposed index utilizes the learning-to-cluster technique, which was developed for image clustering in the field of computer vision, to cluster relevant geo-textual objects and queries together and build an index. Existing learning-to-cluster techniques  [18, 19] rely on partial pairwise similar/dis-similar relationships of images for training, where the pairwise labels are easy to obtain for images. However, such pairwise positive labels for queries and geo-textual objects in our problem are very sparse and high-quality negative labels are absent, which are important for learning-to-cluster. Additionally, when the number of clusters exceeds the number of image classes, the existing learning-to-cluster technique will produce a highly skewed cluster distribution [18], which will hurt the efficiency. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. These informative negative labels combined with ground-truth positive (relevant) labels (e.g., click-through or human annotation, detailed in Section V-A) enable our index to learn both spatial and textual information, thereby clustering relevant objects and queries together while separating irrelevant objects and queries.
@@
  • New Index. We develop a novel machine learning based ANNS index that can speed up embedding based spatial keyword queries. Different from existing ANNS indexes, it extends the learning-to-cluster technique to cluster relevant objects and queries together. To build an effective and efficient index, we propose a novel pseudo-label generation approach. To the best of our knowledge, this is the first index tailored for embedding based spatial keyword queries. Additionally, this is the first geo-textual index to utilize a neural network for retrieval without relying on an explicit tree structure and associated heuristic algorithms.
@@
  • New Relevance Model. We develop a lightweight and effective relevance model. It is featured with a novel learning-based spatial relevance module to learn spatial relevance adaptively, a new weight learning mechanism to determine the weights between textual and spatial relevance, and a dual-encoder module to compute textual relevance.
@@
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets. Experimental results show that LIST outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline. Our implementation is released at https://github.com/Heisenberg-Yin/LIST.
II RELATED WORK
@@
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries [2, 3, 4] are proposed. Among them, Top-k KNN Spatial-Keyword Query (TkQ[2] aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial relevance and text relevance. The earlier work [2] computes the textual relevance with traditional retrieval models like BM25 [10] and TF-IDF [11] and exploits a simple linear function for spatial relevance. However, these traditional methods have limited effectiveness.
@@
To enhance the effectiveness, several deep learning based methods [13, 5, 12] have been developed for query-POI matching, which is essentially spatial keyword query. PALM [13] considers geographic information by using location embedding techniques and combines it with semantic representations for query-POI matching. DrW [5] computes the deep textual relevance on the term level of query and geo-textual objects and uses the attention mechanism to aggregate scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual relevance and spatial relevance. MGeo [12] employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality. MGeo aligns these two modalities into the same latent space and computes relevance scores based on the produced representations. However, these deep learning based methods focus on improving the effectiveness, but ignore the efficiency issue.
Spatio-Textual Indexes. Although various spatio-textual indexes [3, 4, 7, 2, 8, 9, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] have been designed to efficiently answer spatial keyword queries, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for high-dimensional embedding based retrieval models used by deep learning based methods.
@@
Different from previous spatio-textual indexes developed to accelerate the search for top-k objects with traditional relevance models (which is called sparse retrieval), we aim to design an index to accelerate the search for top-k objects with embedding based relevance models (which is called dense retrieval), which is still an open problem for geo-textual data.
@@
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models [34]. The first category of models (e.g., ARC-II [35] and MatchPyramid [36]) calculates the word-level similarity between queries and documents for the final relevance scores. DrW [5] belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM [37]) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of the dual-encoder architecture [38, 39], which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval [40, 41, 42, 43, 44]. For the second category of methods, to support efficient retrieval, the learned embeddings are usually indexed for Approximate Nearest Neighbor Search (ANNS) [34].
@@
ANNS indexes are developed to expedite the top-k search on high-dimensional data. These techniques can be broadly categorized into two types. The first type focuses on searching a subset of database. The representative methods include inverted file index (IVF) based methods [45, 46, 47], hashing-based methods [48, 49, 50, 51], and graph-based methods [52, 17]. These ANNS indexes rely on heuristic algorithms or centroids to find candidates. Specifically, Inverted File Index (IVF) [46] partitions data into clusters using the k-means algorithm and routes queries to clusters based on the query’s distance to the clusters’ centroids. Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index [52] construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings. Hashing-based methods like Locality-Sensitive Hashing (LSH) index [53] generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates. The other type aims to accelerate the search process itself, such as quantization-based methods [46, 54, 55, 56, 54].
Our method belongs to the first category and is orthogonal to the second category. Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatio-textual relevance models. In contrast, existing ANNS indexes are only for textual relevance. (2) Instead of using heuristic algorithms to generate candidates, our proposed index learns from query-object pairs in a data-driven manner and employs neural networks for data partitioning and query routing. Notably, our proposed index utilizes a novel learning-to-cluster technique to group geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
III PROBLEM STATEMENT AND MOTIVATIONS
III-A Preliminary

### 18 Mar 2024 v2

@@ Abstract
With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that considers both spatial and textual relevance, have found many real-life applications. To efficiently handle TkQs, many indexes have been developed, but the effectiveness of TkQ is limited. To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues and there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
@@
To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST). LIST is featured with two novel components. Firstly, we propose a lightweight and effective relevance model that is capable of learning both spatial and textual relevance. Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index to accelerate the top-k search process of the proposed relevance model, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. There exist two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
1 Introduction
@@
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to search for a ‘delicious Italian restaurant’ that is close to the user’s hotel.
@@
Spatial keyword queries have applications in many real-world scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions. To meet diverse user needs, various spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 . Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance. Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (formulated in Section 3.1). According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient . Most of the existing studies congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 on spatial keyword queries focus on improving the efficiency of handling spatial keyword queries. As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
@@
Although many indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited. As discussed earlier, TkQ typically uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue, e.g., synonyms that consist of different tokens may convey the same meanings, which limits their effectiveness (detailed in Section 3.2). For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models. Another limitation arises from using the linear function of distance to calculate spatial relevance. This practice is based on an assumption that the user’s spatial preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and does not hold on our real-life datasets, where the spatial preference for geo-textual objects exhibits a significant (non-linear) decrease with the increase of distance. This mismatch with real-world situations results in limited effectiveness (as to be detailed in Section 3.2).
@@
To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue. To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models. For example, the recently proposed DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings, identifies top-k relevant terms from object description for each query keyword based on the word embeddings, and utilizes an attention mechanism to aggregate the relevance scores of every query keyword as the textual relevance. It also uses the linear function of distance as the spatial relevance. The interactions between word embeddings bring very high query latency. And it relies on interactions between the word embeddings of queries and objects, which cannot be pre-computed until the query-object pairs are given, therefore there is no efficient index designed for it (detailed in Section 3.2).
To address the aforementioned limitations, this work aims to (1) design a lightweight yet effective relevance model, and (2) design an index to expedite the top-k search process of the proposed relevance model.
For the first objective, we propose a lightweight yet effective relevance model for embedding based spatial keyword queries, which capture the semantic meaning of query keywords in one high-dimensional embedding. The relevance model comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module. The dual-encoder module converts query keywords and object descriptions into two separate embeddings and evaluates textual relevance based on the two embeddings. To overcome the limitation of linear distance function, as explained earlier, we propose a new learning-based spatial relevance module, which is capable of learning from real-world datasets to better evaluate spatial relevance. Additionally, a weight learning module is proposed to adaptively learn a weight to balance the textual and spatial relevance.
For the second objective, existing geo-textual indexes are developed based on traditional models, and cannot be used to handle embedding based spatial keyword queries. Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for embedding based spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section 5.2). An intuitive idea is to adapt these ANNS indexes to incorporate the spatial factor. For instance, the IVF index jegouProductQuantizationNearest2011 clusters embeddings using the K-means algorithm and route queries to a small number of close clusters to reduce the search space. However, incorporating spatial factors into the K-means algorithm is challenging as it is hard to set the weight to balance spatial and embedding similarities during index construction. Manually setting the weights is not only laborious but also ineffective, resulting in inferior effectiveness (as to be shown in Section 5.2). It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
@@
To fill this gap, we develop a novel machine learning based ANNS index. To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones. Existing learning-to-cluster studies Hsu16_KCL ; Hsu18_L2C demonstrate that high-quality pairwise similar/dis-similar labels are essential for training. Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse , and high-quality negative labels are absent. Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built on the cluster results. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
Based on the two proposed techniques, namely the lightweight yet effective relevance model and the machine learning based ANNS index, we propose a new solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries (called LIST). LIST employs the proposed relevance model to pre-compute embeddings for geo-textual objects and uses the proposed index to cluster the relevant objects together while separating irrelevant ones. Given embedding based spatial keyword queries, LIST routes queries to a subset of clusters to reduce the search space. Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (detailed in Section 4).
@@
  • New Relevance Model. We develop a lightweight yet effective relevance model. It is featured with a dual-encoder module to convert query keywords and object descriptions into two separate embeddings and compute textual relevance based on the two embeddings, a novel learning-based spatial relevance module to learn from real-word datasets to better evaluate the spatial relevance, and a new weight learning mechanism to determine the weights between textual and spatial relevance.
  • New Index. We develop a novel machine learning based ANNS index that is tailored for embedding based spatial keyword queries and accelerating the top-k search process of the proposed deep relevance model. Different from existing ANNS indexes, it utilizes the learning-to-cluster technique to cluster relevant objects and queries together while separating the irrelevant ones. To build an effective and efficient index, we propose a novel pseudo-label generation approach. To the best of our knowledge, this is the first index designed for embedding based spatial keyword queries and deep relevance model. Additionally, this is the first geo-textual index that employs neural networks for retrieval without relying on an explicit tree structure.
@@
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets. Experimental results show that LIST significantly outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
2 Related Work
@@
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient are proposed. Among them, Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance. However, these traditional methods have limited effectiveness.
@@
To enhance the effectiveness, several deep learning based methods DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 have been developed for query-POI matching, which is essentially spatial keyword query. PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 considers geographic information by using location embedding techniques and combines it with textual word semantic representations for query-POI matching. DrW liu2023effectiveness computes the deep textual relevance on the term level of query keywords and object descriptions. It uses the attention mechanism to aggregate the scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual and spatial relevance. MGeo DBLP:conf/sigir/DingCXHLZX23 employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality. MGeo aligns these two modalities into the same latent space and computes relevance scores based on the produced representations. However, these deep learning based methods focus on improving the effectiveness but ignore the efficiency issue.
Different from previous spatio-textual relevance models that rely on word embeddings and complex models to compute textual relevance, our proposed model employs a dual encoder module to transform query keywords and object descriptions into two embeddings, computes textual relevance based on two embeddings, and we also propose a new learning-based spatial module to learn from real-world datasets to better estimate spatial relevance.
Spatio-Textual Indexes. Various spatio-textual indexes DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:conf/icde/FelipeHR08 ; congEfficientRetrievalTopk2009 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 ; DBLP:conf/icde/ZhangZZL13 ; tao2013fast ; vaid2005spatio ; DBLP:conf/cikm/GobelHNB09 ; DBLP:conf/cikm/ZhouXWGM05 ; DBLP:conf/sigmod/LuLC11 ; DBLP:conf/edbt/ZhangTT13 ; DBLP:conf/icde/ZhangOT10 ; DBLP:conf/icde/ZhangCMTK09 ; DBLP:journals/pacmmod/Sheng0F00C023 ; DBLP:conf/gis/Liu022 have been designed to efficiently answer spatial keyword queries. However, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for accelerating the top-k search process of these deep learning based methods.
@@
Additionally, there exists a learned geo-textual index called WISK DBLP:journals/pacmmod/Sheng0F00C023 , which utilizes query workloads to partition geo-textual data to build a tree-based index and employ reinforcement learning techniques to optimize the index. However, WISK is designed for spatial keyword range queries, which treats query keywords and query region as Boolean filters, only retrieves objects containing all the query keywords within the given query region, and is not designed to support other queries like the TkQ, which uses keywords to compute textual relevance.
@@
Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model, which relies on embeddings to compute textual relevance. It is still an open problem for geo-textual data.
@@
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models guo2020deep . The first category of models (e.g., ARC-II DBLP:conf/nips/HuLLC14 and MatchPyramid DBLP:conf/aaai/PangLGXWC16 ) calculates the word-level similarity between queries and documents for textual relevance. DrW liu2023effectiveness belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM huang2013learning ) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of the dual-encoder architecture DBLP:conf/emnlp/KarpukhinOMLWEC20 ; DBLP:conf/icml/GuuLTPC20 , which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 . For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep . Note that, models in the first category like DrW liu2023effectiveness are not efficient for online computation and retrieval since the complex word interaction function cannot be pre-calculated until we see the input query-object pairs according to guo2020deep .
@@
ANNS indexes are developed to expedite the top-k search on high-dimensional embeddings. These techniques can be broadly categorized into two types. The first type focuses on searching a subset of the database. The representative methods include inverted file index (IVF) based methods DBLP:journals/pami/BabenkoL15 ; jegouProductQuantizationNearest2011 ; wang2020deltapq , hashing-based methods luoSurveyDeepHashing2023 ; liu2014sk ; liu2019lsh ; zheng2020pm , and graph-based methods malkovEfficientRobustApproximate2020 ; DBLP:journals/pvldb/WangXY021 . These ANNS indexes typically rely on heuristic algorithms to find a subset of candidates. Specifically, Inverted File Index (IVF) jegouProductQuantizationNearest2011 partitions data into clusters using the k-means algorithm and routes queries to a subset of close clusters based on the query’s distance to the clusters’ centroids. Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index malkovEfficientRobustApproximate2020 construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings. Hashing-based methods like Locality-Sensitive Hashing (LSH) index DBLP:journals/corr/WangSSJ14 generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates. The other type aims to accelerate the search process itself, such as quantization-based methods jegouProductQuantizationNearest2011 ; DBLP:journals/pami/GeHK014 ; gao2019beyond ; guoAcceleratingLargeScaleInference2020 ; DBLP:journals/pami/GeHK014 .
Our method belongs to the first category and is orthogonal to the second category. Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatial keyword queries. In contrast, existing ANNS indexes are designed for embedding retrieval and do not work well when being adapted to embedding based spatial keyword queries (to be shown in Section 5.2). (2) Existing ANNS techniques are based on heuristic methods. If we want to incorporate spatial factor into the existing ANNS techniques, it is unavoidable to balance the weight between spatial and embedding relevance, which is however challenging. To avoid this, our proposed index employs neural networks to learn to cluster relevant geo-textual objects together while separating the irrelevant ones, and routes embedding based spatial keyword queries to relevant clusters through the neural networks to accelerate the top-k search process. Notably, our proposed index utilizes a novel learning-to-cluster technique to group embedding based geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
3 Problem Statement and Motivations
3.1 Preliminary